The economy's current rate of interest is 10 percent and a firm has NS1000 of owner invested capital is total revenue is NISS000 and the firm's explicit costs are NES3500 . From this we know that this fe's accounting profit is NIS11.500 .


1
Expert's answer
2021-06-02T12:30:04-0400

Solution:

Economic profits = Total revenue – (Explicit costs + Implicit costs)

Total revenue = Accounting profit + explicit costs = 11,500 + 3,500 = 15,000

Accounting profit = Total revenue – explicit costs = 15,000 – 3,500 = 11,500

Implicit costs = "10\%\times 1,000 = 100"

 

Economic profits = 15,000 – (3,500 + 100) = 11,400

 

Economic profits = NIS11,400